Which of the following represents the integral "{\int \frac{1}{x\ln{x}}dx}"


1) "ln(lnx)+C"


2) "ln(xlnx)+C"


3) "lnxlnx+C"


4) None of the above


1
Expert's answer
2021-08-23T15:46:18-0400

"\smallint\frac{1}{xln(x)}dx"


Substitute "u=ln(x)"


Hence "\frac{du}{dx}=\frac{1}{x}," "dx=xdu."


Therefore, = "\smallint\frac{1}{u}du"


This the standard integral;


"=ln(u)+C"


Undo substitution "u=ln(x)"


"=ln(ln(x))+C"

Hence the answer is "ln(lnx)+C"